8x+9y-(2x+y) if x=4 and y=6
Answer by addingup(3677)

Evaluate 8x+9y-(2x+y) where x = 4 and y = 6:
8x+9y-(2x+y) = 8×4+9×6-(2×4+6) Multiply:
32+54-(8+6) Add:
86-14 Subtract:
72 is your answer